Opened 7 years ago

Last modified 6 years ago

#401 assigned flight processing

EU11/03, flight day 142/2011, Jimena

Reported by: knpa Owned by: knpa
Priority: alpha 4 low Milestone: 2011 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: /users/rsg/arsf/arsf_data/2011/flight_data/spain_portugal/EU11_03-2011_142_Jimena

Data arrived from ARSF via network transfer on 24/05/2011

EU11/03 Scientific objective: We wish to measure how these forests change over gradients of topography‐controlled climate and soil properties

EU11/03 Priority: Alpha 4l

PI: David Coomes

Sensors:

Eagle (15/08/2011)
Hawk (15/08/2011)
LiDAR (15/08/2011) and (29/03/12)
Camera (18/07/2011)

Change History (27)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)

comment:2 Changed 6 years ago by anch

Processing RAW camera to tif

comment:3 Changed 6 years ago by anch

Processing IPAS

comment:4 Changed 6 years ago by anch

Produced thumbnails from tifs

comment:5 Changed 6 years ago by anch

IPAS processed.

Produced sol with LIDAR and camera lever arms, and converted one to SBET for hyperspectral processing.

comment:6 Changed 6 years ago by anch

Produced camera ImageEvents1_processed file with IPAS CO

comment:7 Changed 6 years ago by benj

LiDAR processing currently blocked until we get a new boresight flight

comment:8 Changed 6 years ago by benj

Will Simonson obtained additional base station data for SFER base station for this flight date, have downloaded it and put it with the rest of the data for this flight in the repository.

comment:9 Changed 6 years ago by anch

Reprocessed the navigation with new basestation data and got a much better result.

Have generated new SOL files for camera, lidar and SBET for hyperspectral.

Have recreated the camera delivery with reprocessed SOL file and it is ready for delivery check.

Have started processing hyperspectral data with new SBET.

comment:10 Changed 6 years ago by iopa

Camera delivery check ;

  • readme proj code _ to / if tex is available

otherwise all is good and ready to go.

comment:11 Changed 6 years ago by anch

Putting SCT correction on hold until LiDAR can be processed so that it can be verified against an intensity image

comment:12 Changed 6 years ago by emca

  • Description modified (diff)

Camera data sent to PI 18/07/2011

comment:13 Changed 6 years ago by jaho

Obtained intensity image to use for SCT correction. Continuing eagle and hawk processing.

comment:14 Changed 6 years ago by jaho

Hyperspectral data ready for delivery check.

comment:15 Changed 6 years ago by jaho

LIDAR delivery check:

No errors found. Data ready to deliver.

comment:16 Changed 6 years ago by adbe

Delivery check complete.
Data ready to deliver

comment:17 Changed 6 years ago by adbe

  • Description modified (diff)

Data has been delivered
Data on disk 112: -

-142 lidar
-142 hyperspectral
-141 lidar

Delivered to David Coombes on 15/08/2011

comment:18 Changed 6 years ago by adbe

  • Component changed from Processing: general to Archiving
  • Owner set to benj

comment:19 Changed 6 years ago by knpa

  • Milestone set to 2011 data processing completion
  • Owner changed from benj to adbe
  • Status changed from new to assigned

comment:20 Changed 6 years ago by mark1

Hyperspectral hawk bad pixel files and XML files have been re-created and re-delivered to PI (via email). This corrects sample number in bad pixel text files and Eagle IFOV value in XML and extra comments included.

comment:21 Changed 6 years ago by adbe

Hyperspectral sct's

lineeaglehawk
1-0.05-0.05
2-0.02-0.07
3-0.04-0.03
4-0.5-0.10
5-0.120.04
6-0.11-0.11
7-0.01-0.02
8-0.09-0.05
9-0.04-0.04
10-0.03-0.04
11-0.05-0.05
12-0.03-0.05
13-0.05-0.08
14-0.04-0.08
15-0.04-0.03
16-0.03-0.05
17-0.01-0.02
18-0.021.17

comment:22 Changed 6 years ago by emca

Received an email from Will Simonson 09/02/2012 to say the data did not
match up with overlapping lines from day 102 (ticket #395 ).

Vertical offset ~1.2 meters and horizontal shift visible. Reprocessed
the navigation data using a network adjusted temporary base station.
Original data was processed without the correct IBRC file. Data was also
found to have a pitch offset.

Have reprocessed the data and made available to W Simonson via ftp.

comment:23 Changed 6 years ago by knpa

  • Owner changed from adbe to knpa

Beginning archiving.

comment:24 Changed 6 years ago by knpa

No project information xml was included with the delivery.

comment:25 Changed 6 years ago by knpa

  • Description modified (diff)

comment:26 Changed 6 years ago by knpa

  • Description modified (diff)

comment:27 Changed 6 years ago by knpa

Archiving complete. Made available to NEODC.

Note: See TracTickets for help on using tickets.